DFB, Inc., expects earnings this year of $ 5.29 per​ share, and it plans to pay a $ 3.05 dividend to shareholders. DFB will retain $ 2.24 per share of its earnings to reinvest in new projects with an expected return of 15.8 % per year. Suppose DFB will maintain the same dividend payout​ rate, retention​ rate, and return on new investments in the future and will not change its number of outstanding shares.

a. What growth rate of earnings would you forecast for​ DFB?

b. If​ DFB's equity cost of capital is 11.2 %​, what price would you estimate for DFB​ stock?

c. Suppose DFB instead paid a dividend of $ 4.05 per share this year and retained only $ 1.24 per share in earnings. That​ is, it chose to pay a higher dividend instead of reinvesting in as many new projects. If DFB maintains this higher payout rate in the​future, what stock price would you estimate​ now? Should DFB follow this new​ policy?
Answer #1

a). Growth Rate = Retention Rate * ROE

= [$2.24 / $5.29] * 15.8% = 6.69%

b). Price of DFB's Stock = D1 / [r - g]

= $3.05 / [0.112 - 0.0669] = $3.05 / 0.0451 = $67.63

c). Growth Rate = Retention Rate * ROE

= [$1.24 / $5.29] * 15.8% = 3.70%

Price of DFB's Stock = D1 / [r - g]

= $4.05 / [0.112 - 0.0370] = $4.05 / 0.0750 = $54.03

No, projects are positive NPV (return exceeds cost of capital), so don’t raise dividend.